OpenStudy (anonymous):

Jonathan weighs 56.25 pounds. What is his equivalent weight in kilograms? Round off your decimal answer to 2 decimal places. Enter only significant zeros. Help please!

OpenStudy (anonymous):

One pound is 0.453592 kilograms. You would simply multiply 56.25 by 0.453592 which would give you 25.51455 To 2 decimal places: 25.51

OpenStudy (anonymous):

1 kg = 2.2 pounds another way of writing this is\[\frac{1\text{ kg}}{2.2\text{ pounds}}=1\]

OpenStudy (anonymous):

\[56\text{ pounds}\times\frac{1\text{ kg}}{2.2 \text{ pounds}}\]observe that the pounds will cancel